# Obliviate: Efficient Unmemorization for Protecting Intellectual Property in Large Language Models

### v1

@@ 4.2 Obliviate
By combining these two losses, Obliviate effectively unmemorizes specific tokens without compromising the model’s utility. The final loss function that Obliviate optimizes can be expressed as:
@@
where $\lambda_{f}$ and $\lambda_{m}$ are hyperparameters that balance the contributions of the forget loss and the maintain loss, respectively.
@@ 5.1.1 Models
We evaluate Obliviate on four large language models of varying sizes and architectures:
  • LLaMA-3.1 8B (Llama): Base model without instruction tuning (Grattafiori et al., 2024)
  • LLaMA-3.1-instruct 8B (Llama-Inst): Instruction-tuned variant (Grattafiori et al., 2024)
  • Qwen-2.5-7B (Qwen): Base model (Qwen et al., 2025)
  • Yi-1.5 6B (Yi): Base model (01.AI et al., 2024)
@@ 5.1.2 Datasets
We evaluate unmemorization on two different types of datasets:
Synthetic Dataset. We create a synthetic dataset consisting of 10 articles, each approximately 380 words (around 2,000 characters) in length, covering diverse topics. These articles are specifically generated to test the effectiveness of unmemorization on deeply memorized content, where the probability of reconstructing the sequence exceeds 99%. Using synthetic data ensures that the content does not exist in the models’ pretraining data, thereby providing a controlled evaluation environment.
Organic Dataset. We evaluate on excerpts from widely-read literary works, including “Harry Potter and the Philosopher’s Stone”, “Moby Dick”, “Frankenstein”, “Adventures of Huckleberry Finn”, “Alice in Wonderland”, “Sherlock Holmes”, “A Tale of Two Cities”, “Treasure Island”, “Les Misérables”, and “Great Expectations”. These works are particularly relevant as their content has been identified in publicly available models.
Benchmarks. To assess model utility preservation, we evaluate performance on four standard benchmarks, namely HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021b, a), TruthfulQA (Lin et al., 2022), and Winogrande (Sakaguchi et al., 2019).
@@ 5.2 Metrics
Edit Distance (ED). Next, we assess the edit distance using the Levenshtein distance, calculated based on words rather than characters. A lower ED suggests a higher level of memorization by the model. For both LCS and ED, we ignore differences in capitalization.
@@
Figure 2: The average performance of Obliviate in unmemorizing 10 samples across different models, measured using longest Common Subsequence (LCS), edit distance (ED) at the word level, and ROUGE-2 score. 2(a), 2(c) and 2(e) display the results for synthetic data, while 2(b), 2(d) and 2(f) present the results for organic data.
@@ 5.3 Results
We evaluate Obliviate through two main experiments: (1) unmemorizing synthetically memorized data (simulating deeply memorized), and (2) unmemorizing organically memorized data present in the pre-trained models. For each scenario, we assess both the effectiveness of Obliviate memorization and its effect on model utility.
@@
Unmemorizing Synthetic Data. We first finetune the target models on 10 samples using our synthetic dataset until deep memorization is achieved. Intuitively, we finetune the model till its ability to reproduce the complete article given a prefix. We refer to these as “Memorized” models.
@@
Next, we apply Obliviate with a token stride of 5, modifying the token following every set of five tokens, and a modification rate of one token per modification. For the KL-divergence, we consider the top-10 probabilities. We implement early stopping when the probability of the target token falls below a predetermined threshold.
@@
To comprehensively evaluate memorization, we evaluate different testing settings that varies prefix offset positions, i.e., where does the prefix start, from 0 to 200 tokens at 50-token intervals, and prefix lengths ranging from 8 to 20 tokens. We examine both greedy decoding (Greedy) and temperature-based sampling (temperature = 0.6, the default for Llama). For each sample, we report the best performing metrics across all combinations, i.e., the maximum score for Longest Common subsequence (LCS) and ROUGE-2, and the minimum score for Edit Distance (ED). Finally, for the instruct version of the model, we precede the prefix with the following instruction: “Generate the entire rest of this text from the start, continuing until you reach the end: ”.
Figure 2 presents our main results. The LCS metric (2(a)) reveals that memorized models initially reproduce nearly complete articles exceeding 360 words, while Obliviate reduces this to merely 4-5 words— representing a two orders of magnitude reduction in memorization. This finding is further confirmed by the Edit Distance (2(d)), which increases from 5-6 words to over 60 words post-Obliviate. The ROUGE-2 scores (2(e)) similarly demonstrate a significant reduction in verbatim reproduction.
@@
Unmemorizing Organic Data. Our second experiment focuses on organically memorized data present in publicly released models, i.e., without any additional fine-tuning from us. Using the dataset described in Section 5.1.2, we apply the same Obliviate parameters and evaluation methodology as in the synthetic experiments. Results demonstrate comparable effectiveness in removing organic memorization, with LCS (Figure 2(b)) reduced to 2-5 words. Both ED (Figure 2(d)) and ROUGE-2 (Figure 2(f)) metrics show reduction of memorization to negligible levels. Additionally, we present an example in the appendix (Figure 6) illustrating how text generation diverges after unmemorization, resulting in completely different content.
To provide more insight into Obliviate’s effectiveness, we present a sample-level analysis in appendix Figure 7. The per-sample Longest Common Subsequence measurements reveal the consistent significant memorization reduction across all models. While we focus on LCS as an example for the individual samples performance, other metrics (Edit Distance and ROUGE-2) exhibit similar patterns.
@@
Obliviate Effect on Utility. We assess the impact on model utility by comparing Obliviate-unmemorized models against their respective baselines, i.e., the target models on which Obliviate was applied, which are the memorized models for synthetic data and the original models for organic data, across standard benchmarks. Tables 1 and 2 demonstrate the stable performance of Obliviate, with degradation consistently remaining below 1% across all evaluated tasks. This indicates that Obliviate successfully preserves model utility while effectively removing memorized content.
@@
These results demonstrate that Obliviate achieves strong unmemorization performance while maintaining model utility, presenting an efficient approach for selective knowledge unmemorization in large language models. Furthermore, they show that Obliviate effectively generalizes across various model architectures, sizes, and types, including both base and instruction-tuned models. The consistent performance across both synthetic and organic memorization scenarios shows the robustness of our approach across different memorization contexts.
Table 1: Benchmarking Obliviate when Targeting Synthetic Data.
Model Type MMLU TruthfulQA HellaSwag Winogrande
Llama Mem 0.6363 0.4865 0.6084 0.7403
UnMem 0.6349 0.4896 0.6124 0.7514
Llama-Inst Mem 0.6812 0.5387 0.5937 0.734
UnMem 0.6812 0.5367 0.6 0.7411
Yi Mem 0.6237 0.4469 0.5723 0.7214
UnMem 0.6187 0.4520 0.5760 0.7198
Qwen Mem 0.7182 0.5596 0.6032 0.7064
UnMem 0.7169 0.5595 0.6134 0.6851
Table 2: Benchmarking Obliviate when Targeting Organic Data.
Model Type MMLU TruthfulQA HellaSwag Winogrande
Llama Orig 0.6324 0.4518 0.6002 0.7372
unMem 0.6350 0.4501 0.6025 0.7356
Llama-Inst Orig 0.679 0.5402 0.5905 0.738
unMem 0.6772 0.5521 0.5933 0.7403
Yi Orig 0.6242 0.4407 0.5663 0.7198
unMem 0.6276 0.4446 0.5671 0.7198
Qwen Orig 0.7171 0.5637 0.6000 0.7340
unMem 0.7169 0.5737 0.5982 0.7159

### v2

@@ 4.2.2 Maintain Loss
By balancing these objectives, Obliviate unmemorizes selected tokens without compromising model utility. The final objective is
@@
where $\lambda_{f}$ and $\lambda_{m}$ control the trade-off between forgetting and maintenance.
@@ 5.1.1 Models
We evaluate Obliviate on four state-of-the-art large language models (LLMs) that vary in architecture, scale, and training methodology, providing a comprehensive assessment of Obliviate’s robustness across diverse model configurations.
LLaMA-3.1 8B (Base (Llama) and Instruction-Tuned (Llama-Inst))[35]. Developed by Meta AI, LLaMA-3.1 8B is an open-weight decoder-only Transformer model. The instruction-tuned variant, LLaMA-3.1 8B-Instruct, is fine-tuned using supervised instruction-following data and reinforcement learning from human feedback, enhancing its alignment with user intents and safety measures.
Qwen-2.5 7B Base (Qwen)[36]. Qwen-2.5 7B, developed by Alibaba Cloud, is also a decoder-based model trained on an extensive 18 trillion token dataset.
Yi-1.5 6B Base (Yi)[37]. Yi-1.5 6B, developed by 01.AI, is another decoder-based model trained on a 3.1 trillion token corpus encompassing English and Chinese texts.
@@ 5.1.2 Datasets
We evaluate unmemorization on three different types of datasets:
Synthetic Dataset. We create a synthetic dataset consisting of 100 articles, each approximately 380 words (around 2,000 characters) in length, covering diverse topics. These articles are specifically generated to test the effectiveness of unmemorization on deeply memorized content, where the probability of reconstructing the sequence exceeds 99%. Using synthetic data ensures that the content does not exist in the models’ pretraining data, thereby providing a controlled evaluation environment.
Organic Dataset. We evaluate on excerpts from widely-read literary works, including “Moby Dick", “Frankenstein", “Adventures of Huckleberry Finn", “Alice in Wonderland", “Sherlock Holmes", “A Tale of Two Cities", “Treasure Island", “Les Misérables", and “Great Expectations". These works are particularly relevant as their content has been identified in publicly available models.
Unmemorization Benchmarks. Furthermore, we evaluate Obliviate on two different unmemorization/unlearning benchmarks, namely, the Copyright Takedown Evaluation (CoTaEval) [9] and the Machine Unlearning Six-Way Evaluation (MUSE) [10]. CoTaEval introduces a news dataset with 1,000 articles to forget and another 1,000 articles for the retain set. However, we do not use the retain set, as Obliviate does not require any retain set. They also provide an already memorized model, which we directly use for unmemorization. This model is based on the LLaMA2 7B chat version. We use this specific version to directly compare our performance against the other techniques they report.
MUSE [10] introduces another benchmark for comparing unlearning techniques. While Obliviate mainly targets verbatim memorization, we evaluate using MUSE to demonstrate Obliviate’s generalizability. We focus on the News dataset of MUSE, which consists of 889 news articles in the forget set, averaging 1000 tokens each but with samples up to 3600 tokens in length—this is the longest setting we test on. Similar to CoTaEval, MUSE also presents the memorized model (LLaMA-2 7B), which we directly use for Obliviate.
Utility Benchmarks. To assess model utility preservation, we evaluate performance on four standard benchmarks, namely HellaSwag [38], MMLU [39, 40], TruthfulQA [41], and Winogrande [42].
@@ 5.2 Metrics
Edit Distance (ED). Next, we assess the edit distance using the Levenshtein distance, calculated based on words. A lower ED suggests a higher level of memorization by the model. For both LCS and ED, we ignore differences in capitalization.
@@
Figure 2: The worst-case performance of Obliviate across different models, measured using Longest Common Subsequence (LCS) at the word level, Edit Distance (ED) at the word level, and ROUGE-2 score. 2(a), 2(c), and 2(e) display the results for synthetic data, while 2(b), 2(d), and 2(f) present the results for organic data. In each case, we consider the maximum values for LCS and ROUGE-2, and the minimum values for ED to provide a comprehensive evaluation.
Original Paragraph:
the world. It is a way I have of driving off the spleen and regulating the circulation. Whenever I find myself growing grim about the mouth; whenever it is a damp, drizzly November in my soul; whenever I find myself involuntarily pausing before coffin warehouses, and bringing up the rear of every funeral I meet; and especially whenever my hypos get such an upper hand of me, that it requires a strong moral principle to prevent me from deliberately stepping into the street, and methodically knocking people’s
Generated Paragraph After Unmemoriazton:
they can tell you, too, that the life is one of hard work, long hours, and a very uncertain future. The life of a sailor is not for a man who is averse to work. But for those who have the sea in their blood, and [$\cdots$]
TABLE I: Benchmarking Obliviate when Targeting Synthetic and Organic Data.
Model Type MMLU TruthfulQA HellaSwag Winogrande
Organic Synthetic Organic Synthetic Organic Synthetic Organic Synthetic
Llama Mem 0.6324 0.6322 0.4519 0.5110 0.6007 0.6109 0.7388 0.7174
UnMem 0.6337 0.6281 0.4551 0.5044 0.6010 0.6123 0.7419 0.7245
Difference $\uparrow$ 0.21% $\downarrow$ 0.65% $\uparrow$ 0.71% $\downarrow$ 1.29% $\uparrow$ 0.05% $\uparrow$ 0.23% $\uparrow$ 0.42% $\uparrow$ 0.99%
Llama-Inst Mem 0.6790 0.6812 0.5402 0.5387 0.5905 0.5937 0.7380 0.7340
UnMem 0.6773 0.6787 0.5425 0.5348 0.5939 0.5970 0.7356 0.7403
Difference $\downarrow$ 0.25% $\downarrow$ 0.37% $\uparrow$ 0.43% $\downarrow$ 0.72% $\uparrow$ 0.58% $\uparrow$ 0.56% $\downarrow$ 0.33% $\uparrow$ 0.86%
Yi Mem 0.6242 0.6226 0.4409 0.4529 0.5671 0.5800 0.7206 0.7056
UnMem 0.6227 0.6202 0.4409 0.4597 0.5657 0.5819 0.7174 0.7174
Difference $\downarrow$ 0.24% $\downarrow$ 0.39% $\downarrow$ 0.00% $\uparrow$ 1.50% $\downarrow$ 0.25% $\uparrow$ 0.33% $\downarrow$ 0.44% $\uparrow$ 1.67%
Qwen Mem 0.7190 0.7192 0.5638 0.5584 0.6006 0.6037 0.7340 0.7143
UnMem 0.7161 0.7191 0.5690 0.5715 0.5975 0.6176 0.7230 0.6930
Difference $\downarrow$ 0.40% $\downarrow$ 0.01% $\uparrow$ 0.92% $\uparrow$ 2.35% $\downarrow$ 0.52% $\uparrow$ 2.30% $\downarrow$ 1.50% $\downarrow$ 2.98%
(a) Max.
(b) Mean.
Figure 4: Comparison of different copyright takedown methods using CoTaEval benchmarks for both mean and maximum values. Values are normalized, and for ROUGE-1, ROUGE-L, semantic similarity, LCS (word), ACS (word), and Minhash similarity, the values are inverted to ensure that higher values (closer to the boundary) indicate better performance.
Figure 5: Evaluation result on the MUSE benchmark.
(a) Synthetic (LCS $\downarrow$).
(b) Organic (LCS $\downarrow$).
Figure 6: The effect of varying the Top-k probabilities for $\mathcal{L}_{\texttt{forget}}$ and $\mathcal{L}_{\texttt{maintain}}$ using organic and synthetic data.
@@ 5.3 Results
We first evaluate Obliviate through two main experiments: (1) unmemorizing synthetically memorized data (simulating deeply memorized), and (2) unmemorizing organically memorized data present in the pre-trained models. For each scenario, we assess both the effectiveness of Obliviate memorization and its effect on model utility.
@@
Unmemorizing Synthetic Data. We first finetune the target models on 100 samples using our synthetic dataset until deep memorization is achieved, i.e., tokens are memorized up to at least 99% confidence. Intuitively, we finetune the model till its ability to reproduce the complete article given a prefix. We refer to these as “Memorized” models.
@@
Next, we apply Obliviate with a token stride of 4, modifying the token following every set of four tokens, and a modification rate of one token per modification. For the KL-divergence, we consider the top-10 probabilities. We implement early stopping when the forget loss falls below a predetermined threshold, which correlates with target tokens achieving low probabilities.
@@
To comprehensively evaluate memorization, we consider different testing settings that vary prefix offset positions, i.e., where the prefix starts, ranging from 0 to 200 tokens at 50-token intervals, and prefix lengths ranging from 8 to 20 tokens. We examine both greedy decoding (Greedy) and temperature-based sampling (temperature = 0.6, the default for LLaMA). We consider the worst-case scenario in evaluation by considering the absolute maximum for all sentences and inputs for the Longest Common Subsequence (LCS) and ROUGE-2, while considering the minimum Edit Distance (ED). Intuitively, this means that if ALL inputs were 100% forgotten except for one, this evaluation method would completely indicate it as a failure. For completeness, we also show the mean results; however, we believe the maximum/minimum metrics are the best indicators of how successful an unmemorization technique is. Finally, for the instruct version of the model, we precede the prefix with the following instruction: “Generate the entire rest of this text from the start, continuing until you reach the end: ”.
Figure 2 demonstrates the efficacy of our approach across multiple evaluation metrics and model architectures. The Longest Common Subsequence (LCS) analysis reveals a dramatic reduction in memorization: while baseline models reproduce substantial portions of training data—with memorized sequences, for instance, Qwen (2(a)) generates LCS of size 412 words, which is reduced to 4-5 words after applying Obliviate, achieving over two orders of magnitude improvement. Similar patterns are observed across different models and datasets. This is also demonstrated by the Edit Distance results, which significantly increase after applying Obliviate for all models and datasets, and by the ROUGE-2 scores, which provide additional confirmation, demonstrating the significant unmemorization achieved by Obliviate.
@@
Unmemorizing Organic Data. Our second experiment focuses on organically memorized data present in publicly released models, i.e., without any additional fine-tuning from us. Using the dataset described in Section 5.1.2, we apply the same Obliviate parameters and evaluation methodology as in the synthetic experiments. Results demonstrate comparable effectiveness in removing organic memorization, with LCS (Figure 2(b)) reduced to 2-5 words. Both ED (Figure 2(d)) and ROUGE-2 (Figure 2(f)) metrics show reduction of memorization to negligible levels. Additionally, we present an example in Figure 3 illustrating how text generation diverges after unmemorization, resulting in completely different content.
@@
Obliviate Effect on Utility. We assess the impact on model utility by comparing Obliviate-unmemorized models against their respective original models across standard benchmarks. Table I demonstrates the stability of Obliviate, with performance degradation consistently remaining minimal across all evaluated tasks and model architectures.
[…]
In general, across all 32 evaluation scenarios (4 models × 4 benchmarks × 2 datasets), the maximum degradation observed is 2.98% on Winogrande for Qwen with synthetic data, while the majority of cases show degradations well below 1% and some an improvement. This consistent performance across diverse model architectures—from base models (Llama, Yi, Qwen) to instruction-tuned variants (Llama-Inst)—demonstrates that Obliviate successfully preserves model utility while effectively removing memorized content.
@@
These results establish that Obliviate achieves strong unmemorization performance without sacrificing model capabilities, presenting an efficient approach for selective knowledge removal in large language models. The consistent utility preservation across various model sizes, architectures, and both synthetic and organic memorization scenarios show the robustness and practical applicability of our approach across different deployment contexts.
Method ROUGE1 ROUGEL Sem Sim LCS ACS LD Minhash MMLU MT-Bench Blocked F1 In-Domain F1
Sys Prompt 0.839 0.801 0.930 119 124 1107 0.812 0.340 0.330 0.334 0.330
PO 0.579 0.515 0.868 37 65 1230 0.445 0.331 0.240 0.283 0.247
Grad Diff 0.362 0.191 0.842 14 14 1039 0.289 0.158 0.150 0.169 0.162
Grad Ascent 0.477 0.270 0.876 21 33 1107 0.367 0.279 0.330 0.269 0.258
KL 0.409 0.229 0.807 14 20 1035 0.336 0.176 0.150 0.169 0.159
Obliviate 0.470 0.280 0.877 15 40 1140 0.427 0.357 0.467 0.366 0.370
TABLE II: CoTaEval results.
5.4 Generalizability and Comparison
Next, we demonstrate the generalizability of Obliviate by evaluating it against two benchmarks: CoTaEval [9] and MUSE [10].
CoTaEval. We begin with CoTaEval, which encompasses multiple copyright-enforcement mechanisms, including both training-time approaches (similar to Obliviate) and test-time approaches. In this work, we focus primarily on the comparison with training-time techniques that are most similar to Obliviate, namely Gradient Ascent (Grad Ascent)[11], Gradient Difference (Grad Diff)[12], KL Minimization (KL)[13], and Preference Optimization (PO)[14]. We also include a simple unmemorization baseline that explicitly states the forget request in the system prompt (Sys Prompt). All methods are applied to a Llama-2 7B chat model that has been memorized on the NewsQA dataset (it is provided by CoTaEval). The NewsQA forget set consists of 1,000 inputs to forget, which further tests the robustness of each method when scaling to larger datasets.
CoTaEval evaluates forgetting-effectiveness using the following metrics:
  • ROUGE-1 / ROUGE-L[43]: unigram and longest-common-subsequence overlap between the generated output and the copyrighted/target content to forget;
  • LCS (character) / LCS (word): longest common subsequence measured at the character and word levels;
  • ACS (word): accumulated common subsequences in words;
  • Levenshtein Distance[44]: the edit distance between the generated output and the blocklisted content;
  • Semantic Similarity: cosine similarity between the generated output and the copyrighted/target content to forget;
  • MinHash Similarity [45]: the Jaccard-approximation similarity via MinHash sketches.
To assess overall model utility, CoTaEval additionally measures: MMLU (Multi-Task Language Understanding) accuracy and MT-Bench [46] scores.
Finally, to quantify in-domain knowledge retention for inputs not in the forget set, CoTaEval reports word-level F1 on (i) blocklisted content and (ii) in-domain content that should not be forgotten.
4(a) shows the comparison of all unmemorization techniques on CoTaEval using maximum metric values, and 4(b) shows the same comparison using mean metric values (the utility metrics remain identical in both figures). Both figures are normalized and adjusted—so that for ROUGE-1, ROUGE-L, Semantic Similarity (Semantic Sim), LCS(word), ACS(word) and MinHash Similarity, lower raw scores are inverted (since lower values are better for these metrics) to yield higher normalized values, i.e., closer to the boundary is always best. Grad Diff achieves the best performance with respect to ROUGE-1 (0.362 vs. 0.470 for Obliviate), ROUGE-L (0.191 vs. 0.280 for Obliviate), LCS(word) (14 vs. 15 for Obliviate), ACS(word) (14 vs. 40) and MinHash (0.289 vs. 0.427), while KL Minimization outperforms Obliviate only on Semantic Similarity (raw 0.807 vs. 0.877 ) and PO outperforms Obliviate only on Levenshtein Distance (raw 1,230 vs. 1,140). In contrast, Obliviate dominates all utility metrics—MMLU (0.357), MT-Bench (4.67), Blocklisted F1 (0.366) and In-Domain F1 (0.370)—demonstrating the best overall utility. This result shows that Obliviate strikes a more balanced trade-off by combining near–state-of-the-art forgetting performance with significantly higher utility. We present the concrete numbers for the maximum evaluation of CoTaEval in Table II.
MUSE. Second, we evaluate MUSE on the News dataset (distinct from the dataset used in CoTaEval). This dataset comprises 889 samples to forget, averaging approximately 1000 tokens in length with samples as long as 3600 tokens. Due to memory constraints, we split the inputs into segments of at most 512 tokens, resulting in an increased number of forget-inputs. The MUSE inputs are heavily formatted with punctuations, which had an effect on the performance of Obliviate. To address this, we simply modified the stride selection to ignore punctuations, i.e., if punctuations are selected as target tokens to unmemorize. Instead, the Obliviate would search for the next token, first by going to the previous tokens, and if not found, then searching the next tokens. After selecting a new target token, the stride would start from this new position and continue normally.
The MUSE benchmark also employs the ROUGE-L metric in two distinct modes: No Verbatim Memorization (VerbMem ROUGE), which ensures that the model does not reproduce the exact input, thus measuring memorization; and No Knowledge Memorization (KnowMem ROUGE), which requires the model to erase all knowledge from the input, testing complete unlearning—a capability Obliviate does not aim to provide. Additionally, MUSE uses Extraction Strength to quantify how closely the model’s predictions align with ground-truth labels. To assess performance on data that should not be forgotten, MUSE also computes ROUGE-L on the “retain” subset (Retain KnowMem ROUGE). Beyond MUSE’s utility evaluation, we evaluate all techniques across four additional benchmarks: MMLU, TruthfulQA, WinoGrande, and HellaSwag. We do not use MT-Bench here as it is costly to run since it uses GPT-4 as a judge.
MUSE considers gradient-based methods analogous to those in CoTaEval (Grad Ascent and Grad Difference), as well as negative preference optimization (NPO) [15], SimNPO [16], and the original memorized model (Target) as the unlearning techniques.
We follow the same normalization procedure from CoTaEval. We invert VerbMem ROUGE, KnowMem ROUGE, and Extraction Strength before plotting, since lower scores indicate better performance for these metrics. The results, shown inFigure 5, reveal that, similar to CoTaEval, Grad Difference most closely approaches Obliviate’s overall performance; however, Obliviate still achieves higher utility, nearly matching the Target model. This demonstrates that Obliviate can effectively unmemorize the target inputs while maintaining strong utility. Finally, it is important to recap that unlike other methodologies, Obliviate does not require a retain dataset, reducing the number of assumptions needed for execution.
(a) Synthetic (LCS $\downarrow$)
(b) Organic (LCS $\downarrow$)
(c) Benchmark Llama Synthetic
(d) Benchmark Llama Organic
(e) Benchmark Qwen Synthetic
(f) Benchmark Qwen Organic
Figure 7: Performance of Obliviate with varying stride lengths. 7(a) and 7(b) illustrate the LCS on synthetic and organic data, respectively. 7(c) and 7(d) present benchmark results for synthetic and organic data on the Llama model, while 7(e) and 7(f) present results for the Qwen model.